Use the matrix method (together with elementary row transformations) to solve the following:


2x −y +3z = 2


x +2y −z = 4


−4x +5y +z = 10


1
Expert's answer
2022-06-15T14:56:35-0400
"A=\begin{pmatrix}
 2 & -1 & 3 \\
 1 & 2 & -1\\
 -4 & 5 & 1 \\
\end{pmatrix}, X=\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}, B=\begin{pmatrix}
 2\\
 4\\
10
\end{pmatrix}"

"AX=B"

"A^{-1}AX=A^{-1}B"

"X=A^{-1}B"

Augment the matrix with the identity matrix:


"\begin{pmatrix}
 2 & -1 & 3 && 1& 0 & 0 \\
 1 & 2 & -1 && 0 & 1 & 0\\
 -4 & 5 & 1 && 0 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
 1 & -1/2 & 3/2 && 1/2& 0 & 0 \\
 1 & 2 & -1 && 0 & 1 & 0\\
 -4 & 5 & 1 && 0 & 0 & 1 \\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & -1/2 & 3/2 && 1/2& 0 & 0 \\
 0 & 5/2 & -5/2 && -1/2 & 1 & 0\\
 -4 & 5 & 1 && 0 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3+4R_1"


"\begin{pmatrix}
 1 & -1/2 & 3/2 && 1/2& 0 & 0 \\
 0 & 5/2 & -5/2 && -1/2 & 1 & 0\\
 0 & 3 & 7 && 2 & 0 & 1 \\
\end{pmatrix}"

"R_2=2R_2\/5"


"\begin{pmatrix}
 1 & -1/2 & 3/2 && 1/2& 0 & 0 \\
 0 & 1 & -1 && -1/5 & 2/5 & 0\\
 0 & 3 & 7 && 2 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1+R_2\/2"


"\begin{pmatrix}
 1 & 0 & 1 && 2/5 & 1/5 & 0 \\
 0 & 1 & -1 && -1/5 & 2/5 & 0\\
 0 & 3 & 7 && 2 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3-3R_2"


"\begin{pmatrix}
 1 & 0 & 1 && 2/5 & 1/5 & 0 \\
 0 & 1 & -1 && -1/5 & 2/5 & 0\\
 0 & 0 & 10 && 13/5 & -6/5 & 1 \\
\end{pmatrix}"

"R_3=R_3\/10"


"\begin{pmatrix}
 1 & 0 & 1 && 2/5 & 1/5 & 0 \\
 0 & 1 & -1 && -1/5 & 2/5 & 0\\
 0 & 0 & 1 && 13/50 & -6/50 & 1/10 \\
\end{pmatrix}"

"R_1=R_1-R_3"


"\begin{pmatrix}
 1 & 0 & 0 && 7/50 & 8/25 & -1/10 \\
 0 & 1 & -1 && -1/5 & 2/5 & 0\\
 0 & 0 & 1 && 13/50 & -6/50 & 1/10 \\
\end{pmatrix}"

"R_2=R_2+R_3"


"\begin{pmatrix}
 1 & 0 & 0 && 7/50 & 8/25 & -1/10 \\
 0 & 1 & 0 && 3/50 & 7/25 & 1/10\\
 0 & 0 & 1 && 13/50 & -6/50 & 1/10 \\
\end{pmatrix}"

We are done. On the left is the identity matrix. On the right is the inverse matrix.

"A^{-1}B=\begin{pmatrix}
 0.14 & 0.32 & -0.1 \\
 0.06 & 0.28 & 0.1 \\
 0.26 & -0.12 & 0.1 \\
\end{pmatrix}\begin{pmatrix}
 2\\
 4\\
10
\end{pmatrix}"

"=\begin{pmatrix}
 0.28+1.28-1\\
 0.12+1.12+1\\
0.52-0.48+1
\end{pmatrix}=\begin{pmatrix}
 0.56\\
 2.24\\
1.04
\end{pmatrix}"

"(x,y,z)=(0.56, 2.24, 1.04)"